Qd=25000-2P

Qs=1000+1P

Calculate the equilibrium of price and quantity of a housing unit


1
Expert's answer
2021-05-10T16:19:54-0400

At equilibrium "Qd=Qs"

"25000-2p=1000+1p"

"25000-1000=1p+2p"

"24000=3p"

"p=8000"


We can put the value of P in any of the equation (demand or supply).

Putting the value on P in demand equation;

"Qd = 25000-2P"  

   "= 25000 - 2(8000)"

   "= 25000 - 16000"

"= 9000"

Hence, the equilibrium of price is 8000 and quantity of a housing unit is 9000.